OpenStudy (acespeedfighter):

Evaluate the expression 7x2y, when x = 5 and y = 3. (3 points) 210 315 525 728 3. (06.04) Evaluate the expression 3.14(a2 + ab) when a = 4 and b = 3. (Input decimals only, such as 12.71, as the answer.) (4 points)

OpenStudy (acespeedfighter):

2 different question

OpenStudy (acespeedfighter):

@ganeshie8

OpenStudy (sachintha):

You need to substitute \(x\) and \(y\) with their respective values. So expression 1 becomes, \(7x2y=7\times5\times2\times3\)

OpenStudy (sachintha):

Do the same thing to the second expression.

OpenStudy (acespeedfighter):

hmmm still dont get how you got the oh wait!!!

OpenStudy (acespeedfighter):

4x3x2 right?

OpenStudy (acespeedfighter):

or 3.14x2x4x3x12.71

OpenStudy (sachintha):

Hmm there is one problem. Is this the correct expression, \(7x^2y\) and \(3.14(a^2+ab)\) ?

OpenStudy (acespeedfighter):

i already did number 2 its 210

OpenStudy (acespeedfighter):

Im looking at 3.

OpenStudy (acespeedfighter):

@MARC_D Cause why not

OpenStudy (sachintha):

Wait, check the expressions that I typed with those in the book. If it's \(7x^2y\) the answer for the first question is different.

OpenStudy (acespeedfighter):

k

OpenStudy (acespeedfighter):

I think it is

OpenStudy (acespeedfighter):

@Sachintha it is

OpenStudy (sachintha):

Which one do you mean by 'it is'? Is it \(7x2y\) or \(7x^2y\) ?

OpenStudy (acespeedfighter):

The 2nd ONE

OpenStudy (sachintha):

Ohh then it should be as follows, \(7x^2y=7\times(x\times x)\times y\)

OpenStudy (acespeedfighter):

oh ok thx

OpenStudy (sachintha):

\(x^2 = x \times x\) I hope you can do the second one now. :)

OpenStudy (acespeedfighter):

Yea i can thx :D

OpenStudy (sachintha):

Great! :)